In the last election, candidate A received twice as many votes as candidate B. Candidate C received 15,000 fewer votes than candidate B. If a total of 109,000 votes were cast, hown many votes did candidate A receive?
Answer by checkley77(12569)

A+B+C=109,000
A=2B
C=B-15,000
2B+B+B-15,000=109,000
4B=109,000+15,000
4B=124,000
B=124,000/4
B=31,000 ARE THE B VOTES
A=2*31,000=62,000 ARE THE A VOTES
C=31,000-15,000=16,000 ARE THE C VOTES
PROOF:
62,000+31,000+16,000=109,000
109,000=109,000